OpenStudy (anonymous):

a formula about a square defines a quantity P to be equal to the sum of all equal four sides (a,b,c and, d ) of a square. solve the formula for when P=28

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Okay... a,b,c, and d are the sides of a *square* right?

OpenStudy (anonymous):

yesh :3

terenzreignz (terenzreignz):

Well, we know that the sides of a square are all equal, right? So a=b=c=d Catch me so far?

OpenStudy (anonymous):

yesh i ended up with 7 but idk how i did lol

terenzreignz (terenzreignz):

You did? Well, see, you have P = a+b+c+d right?

terenzreignz (terenzreignz):

Since P is the sum of all four sides...

OpenStudy (anonymous):

yes and P=28

terenzreignz (terenzreignz):

Yup. But since a,b,c, and d are all equal, you can just have it P = a+a+a+a = 4a right?

OpenStudy (anonymous):

yesh

terenzreignz (terenzreignz):

Well, you have P = 28 P = 4a So... 28 = 4a Solving for a ?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

right???

terenzreignz (terenzreignz):

Yes. And since a = 7 same goes for b,c, and d :P

OpenStudy (anonymous):

yay thank you